OpenStudy (anonymous):

if you have a box and the lenghth is 12cm and the width is 5cm and the width is increasing at 2 cm/sec and the length is decreasing at 2 cm/sec what is the DA/dt for the area perimeter and diagonal

OpenStudy (anonymous):

A=l*b dA/dt = d(l*b)/dt=dl/dt *b + l*db/dt= -2b+2l

OpenStudy (anonymous):

i honestly have no idea which variable is what

OpenStudy (anonymous):

l is length and b is width

OpenStudy (anonymous):

okay so area is 14cm/sec and perimiter is 0 but the diagonal?

OpenStudy (anonymous):

diagonal^2 = l^2 +b^2 x^2=l^2+b^2 2x dx/dt = 2l dl/dt + 2b db/dt

OpenStudy (anonymous):

where x= diagonal

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

okay so pythagerous

OpenStudy (anonymous):

pythagorus*

OpenStudy (anonymous):

can the answer be negative? amd do you know the answer?

OpenStudy (anonymous):

yes the answer can be negative

OpenStudy (anonymous):

x^2=l^2+b^2 x^2=12^2 +5^2 x=13 Now, 2x dx/dt = 2l dl/dt + 2b db/dt 2*13 dx/dt = 2*12*(-2) +2*5*2

OpenStudy (anonymous):

U need to find dx/dt which looks like is negative